Coronal hard X-ray and γ-ray sources

S. Krucker and the ISSI CornSource Team

Space Sciences Lab, UC Berkeley

Abstract. These are notes meant to capture some of the historical development (slow until recently) of this subject, including whatever I think is “known” already from RHESSI data, but it’s a rapidly developing field! - HH

The solar corona supports many forms of activity that in the past have been best observed at radio wavelengths. Over the years, however, many observations of hard X-ray and γ-ray emission have also been reported. Most of these observations have not had the benefit of imaging, so that in many cases they have provided limited physical information. Now RHESSI^{1} has given us a systematic view of coronal sources that combines high spatial and spectral resolution. This paper surveys the RHESSI observations in the context of prior knowledge, emphasizing new phenomena observed in the corona not directly associated with the impulsive phase. Despite the low density and hence low bremsstrahlung efficiency of the corona, we now detect coronal hard X-ray emission from sources in all phases of solar flares.

Keywords: Corona, flares, particle acceleration, X-rays

Table of Contents

1 Introduction 2 2 Basic physics 4 3 Gradual late-phase sources 5 4 The impulsive phase: which ones are the Masuda sources? 6 5 Gradual early-phase sources 8 6 Type III bursts 8 7 Parallel motions 10 8 Footpoint-to-looptop ratio 10 9 Collapsing traps 11 10 Conclusions 11

1 The Reuven Ramaty High-Energy Solar Spectroscopic Imager; see Lin et al. (2002)

\begin{matrix}c \\ ◯ \end{matrix} 2007 Kluwer Academic Publishers. Printed in the Netherlands.

1. Introduction

Energetic electrons, accelerated outside the thermal background distribution, underpin a wide variety of coronal radio sources. One might expect that these “tail population” electrons would also emit detectable levels of hard X-radiation via bremsstrahlung, which would indeed provide a more direct and less model-dependent means of characterizing them. However the hard X-ray/soft γ-ray energy range of a few keV to 10 MeV presents major challenges for sensitive instrumentation. In particular focusing optics become difficult, so that all information to date has come from the relatively high-background imaging techniques based upon image modulation (e.g., Prince et al., 1988). The RHESSI spacecraft, launched in February 2002, is now nevertheless providing a much superior observational view of this difficult domain. In this paper we describe several new types of coronal hard X-ray emission that arguably require non-thermal distribution functions accelerated away from the background Maxwellian plasma.

Thus far the coronal hard X-ray observations have been published essentially event-by-event, although Cliver et al. (1986) and Bai and Sturrock (1989) have reviewed late-phase “extended events,” which Kiplinger (1995) then showed to be related to the production of “solar cosmic rays” (SEPs, solar energetic particles appearing in the heliosphere). These results were based mainly upon non-imaging observations; since this era we have had four hard X-ray imaging experiments (on the Solar Maximum Mission, Hinotori, Yohkoh, and now the RHESSI spacecraft, the only source of hard X-ray and γ-ray imaging for the foreseeable future). These data have allowed us to go beyond the primitive observational technique of limb occultation, and to begin to observe hard X-rays against the disk of the Sun as is common at soft X-ray, EUV, and radio wavelengths. The limb-occultation technique, however, led to our first empirical knowledge of coronal hard X-ray emission, and the stereoscopic comparison of simultaneous data with heliospheric spacecraft such as Ulysses and Pioneer Venus Orbiter led to the early recognition of compact footpoint hard X-ray sources as a characteristic of the impulsive phase (Hoyng et al., 1976; Kane et al., 1979).

The Yohkoh hard X-ray imaging observations gave us the discovery of the celebrated Masuda et al. (1994) “above-the-loop-top” source, an unexpected coronal structure appearing in the impulsive phase of a flare just at the limb. The coronal source was apparently not directly connected with the soft X-ray magnetic loops that characterize soft X-ray images of solar flares. The RHESSI data do not show such an extreme version of this type, but do reveal many other examples of coronal hard X-ray sources as reviewed in these notes. These include what appear to be fundamentally new classes of events, notably hard X-rays appearing prior to the impulsive phase as in the γ-ray flare event of 2002 July 23 (Lin, 2002). We discuss these different types of events, as summarized in Table I, in approximately chronological order; roughly the latter half of the table consists of original RHESSI discoveries. The listing includes the so-called “superhot” sources discovered by Lin et al. (1981) for completeness even though they are apparently thermal in nature; the RHESSI data are now revealing a wide variety of intermediate cases because of its unique access to the 3-12 keV band. We also summarize the basic physics of coronal hard X-ray emission in Section 2. Although it is certainly premature to classify the small number of events that have been reported, several morphologies appear and it is likely that different physics is involved in different events. Our classifications are probably not all mutually exclusive, nor comprehensive. The radio data guarantee that other (albeit faint) sources must occur, and in Section 6 we discuss one of the likeliest possibilities: the direct hard X-ray detection of electrons responsible for type III radio bursts

2. Basic physics

Coronal emission in hard X-rays results from the bremsstrahlung emission of weakly relativistic (> 10 keV) electrons on ambient coronal plasma (more exotic mechanisms are possible but not likely; see Korchak, 1967). We do not have much detailed information about the distribution functions of the particles, nor of the physical state of the plasma in the corona, especially during disturbed conditions. The coronal hard X-ray sources discussed in this paper include (for completeness) those probably only representing the tail of essentially thermal distributions. These “superhot” sources (Lin et al., 1981) presumably represent strongly heated plasmas in which one may expect distortions from a thermal velocity distribution. The RHESSI observations, with good spectral and spatial resolution, may help in understanding these distortions, which may be represented in the coronal hard X-ray sources listed in Table I.

For ordinary flare hard X-ray sources one makes the “thick target” assumption in which the electrons actually stop (Brown, 1971; Hudson, 1972) while radiating. This simplifies the discussion by removing the ambient density as an explicit physical parameter. For coronal sources instead we must think of ordinary thin-target bremsstrahlung.

Table I. Prototypical coronal hard X-ray sources

Description  Prototype event Reference
Late phase, hard spectrum  30 March 1969  Frost & Dennis, 1971
Late phase “superhot” spectrum  21 June 1981  Lin et al. 1981
Impulsive phase “above the loop top” 13 January 1991 Masuda et al. 1994
Fast hard X-ray ejection  18 April 2001  Hudson et al. (2001)
Early-phase hard spectrum  23 July 2002  Lin et al., 2003
Coronal thick-target source  14 April 2002  Veronig & Brown 2004
Early-phase “superhot” spectrum  3 November 2003 Sui et al., in prep?
Coronal γ-ray emission  20 January 2005 Krucker et al., in prep

In the Appendix we present simple formulas convenient for representing the physical parameters of the energetic electrons these coronal hard X-ray sources. The structures in which these electrons find themeselves are almost totally unknown except for the generic concept of the magnetic loops in which they may be trapped by the mirror force. Standard scenarios envisioned for eruptive flares involve a drastic restructuring of these loops, and the presence of solar energetic particles (SEPs) in interplanetary space may mean that some of these same electrons may actually escape from their coronal traps (Kontar & Krucker, in preparation). The stable trapping of electrons is problematic because of the possibility of loss-cone instabilities, but we know little about the 3D structure of the field and can only guess at the mirror ratio (ratio of maximum to minimum field) for a given loop feature. The plasma itself in these features may be distributed in a non-uniform manner, as discussed below and as clearly seen in images of flare ejecta.

3. Gradual late-phase sources

The event of 30 March 1969 (Frost and Dennis, 1971) occurred in an active region known circumstantially to have been some distance behind the solar limb (Smerd & Weiss, 1971), so that the hard X-rays visible from the Earth-orbiting OSO-5 spacecraft probably originated from relatively high in the corona. Two other examples of quite similar events were reported with OSO-7 data (Hudson, 1978; Hudson et al., 1982); such events are characterized by flat hard X-ray spectra (power-law index γ ∼ 2), gradual time profiles, low microwave peak frequencies, anomalously weak soft X-ray emission, and association with coronal

Figure 1. RHESSI non-imaging hard X-ray observations of an extended hard X-ray event, as analyzed by Grigis and Benz (2004), showing the joint variation of spectral flux at 35 keV and spectral index γ. Note the steady evolution from spiky bursts exhibiting soft-hard-soft morphology, upon which a more gradual hardening component can be discerned.

radio bursts. RHESSI has now observed sources in this category (see below) but the ten years of Yohkoh did not produce a clear example from HXT^{1}; we attribute this to the rarity of prototypical events and to the low dynamic range of modulation-based imaging instruments.

Nevertheless a related pattern appears in the so-called “soft-hard-harder” spectral evolution of many long-duration hard X-ray events (Cliver et al., 1986; Kiplinger, 1995). This spectral pattern differs from the otherwise ubiquitous “soft-hard-soft” pattern associated with the impulsive phase (Parks and Winckler, 1969; Hudson and Fárník, 2002; Grigis and Benz, 2004). From a non-imaging perspective, the hard X-ray spectrum of such a source consists of a gradual, continuously hardening component plus a serie.s of spikes with soft-hard-soft evolution. Often these spikes become more gradual as the event develops, as illustrated in Figure 1 (and as seen in the earliest, non-imaging observations, from the TD-1A spacecraft; see Hoyng et al., 1976).

The RHESSI imaging observations have added a new feature for the interpretation of these late-phase sources. Because of their long duration, a thin-target explanation might be imagined, such that the hard X-ray emission would come predominantly from the corona. However it now appears (Krucker. in preparation 2007) that these late sources

1 Yohkoh’s Hard X-ray Telescope; see Kosugi et al., 1992)

Figure 2. Footpoint sources in the late phase of one of the January 2005 flares.

emit hard X-radiation mainly from footpoints, as with ordinary flare loops. Figure 2 illustrates this for one of the flares of January 2005. The footpoint nature of these late-phase events had already been pointed out by (Qiu et al., 2004).

4. The impulsive phase: which ones are the Masuda sources?

The Yohkoh observations of what is now universally termed the “Masuda flare” showed the presence of high-energy electrons in the corona during the impulsive phase; it is possible that the pioneering stereoscopic observations of Kane et al., 1979) spotted a similar phenomenon. Yohkoh directly imaged the hard X-rays from the 13 January 1992 Masuda flare, revealing it to lie above the loop tops of what appeared to be a classical loop flare with bright footpoints. This source occurred during the impulsive phase of the flare, in a region of low density (i.e., with no bright soft X-ray emission nor apparent magnetic connection to the footpoints), so it immediately raised questions about the bremsstrahlung efficiency (Wheatland and Melrose, 1995; Hudson and Ryan, 1995; Fletcher and Martens, 1998). At the same time the morphology matched the cusp geometry taken to imply the presence of an essentially vertical current sheet above the flaring loops. The evident presence of non-thermal particles in this geometry then led to much discussion of the classical large-scale magnetic reconnection models of solar flares, as pointed out by Masuda et al. (1994).

The Masuda observation proved difficult to reproduce, but extensive further analysis of the original 1992 flare continued (e.g., Alexander and Metcalf, 2002; Aschwanden et al., 1999; Metcalf and Alexander, 1999; Petrosian et al., 2002). The novel features of the original Masuda flare were its location above the soft X-ray looptop during the impulsive phase, and a non-thermal spectral signature. There is some confusion in the literature between this source and the more common thermal loop-top sources, which have their own interesting properties. This confusion has been amplified by the unusual characteristics of these ordinary loop-top sources; they generally have a bright soft Xray feature at the apparent loop top (Acton et al., 1992), which can might be interpreted as an anomalous overpressure because the soft X-ray emissivity typically increases monotonically with pressure, ie as a function of the product nT (Kahler, 1976). However any such overpressure should drive flows that eliminate it promptly. Alternatively, the loop-top brightening might be evidence of line-of-sight effects in a cusp structure, which however would suggest a high plasma beta.

A further characteristic of the loop-top sources is the clear existence of a gradient in temperature towards greater heights in a flare loop system. Such a gradient is naturally explained by the standard 2D reconnection model, in the sense that newly-reconnected loops appear always at greater heights, and since there has been no time for evaporation to occur and establish a quasi-static equilibrium loop structure, the plasma temperature should be at least instantaneously higher at the loop tops. Such an effect is often dramatically evident in the TRACE 195Å images. This channel has a spectral passband with two maxima, one showing hotter loops (Fe xxi response) at higher altitudes, and the other showing loops after cooling to ordinary coronal temperatures (the nominal Fe xi, xii passband; see, e.g., Warren et al., 1999).

5. Gradual early-phase sources

The RHESSI observations have revealed what appears to be a new class (or classes) of coronal hard X-ray emission, namely sources that precede the impulsive phase. The prototype of this class is the event of 23 July 2002 (Lin et al., 2003; Asai et al., 2004; Asai et al. 2006), as illustrated in Figure 3. Lin et al. (2003) pointed out that the hard X-ray spectrum in this phase of the event appears almost continuous to low energies, but with the characteristic Fe emission feature at 6.7 keV establishing that some background thermal plasma exists in the source. Another example in this category is the RHESSI event of 3 November 2003 01:09. We speculate that HXT did not observe

Figure 3. RHESSI (left, 25-40 keV and 40-60 keV contours in yellow and blue; Nobeyama 34 GHz contours in red; right, RHESSI 12-25 keV contours in white and Nobeyama flux 34/17 GHz flux ratio interpreted as a power-law index). The RHESSI emission at (-885, -235) comes from the coronal source discussed in the text, and the Nobeyama spectral data on the right confirm its nonthermal nature (From Asai et al., 2004). The background image on the left is TRACE 195Å.

such events readily because of its “flare mode” operation, in which hard X-ray spectra were not telemetered until the hard X-ray flux reached a given threshold, which usually corresponded to the onset of the main impulsive phase. Thus an interesting source prior to the impulsive phase would systematically not have been properly observed (but see Fárník et al., 2003, for an early-phase observation resulting from an unprogrammed mode change).

6. Type III bursts

The direct detection of non-thermal electrons via RHESSI imaging spectroscopy remains in its early stages of development. One wellanticipated type of source would be the hard X-ray counterpart of the type III radio burst, interpreted since the earliest times (e.g. Wild et al., 1963) as the outward motion of streams of non-thermal electrons, as directly confirmed in interplanetary space (e.g., Lin, 1970). Type III bursts often occur during solar flares, especially during the impulsive phase. but also often in the absence of major flares (Kane et al., 1974). At such times soft X-ray observations tend to show the presence of jets, invariably with a microflare near the foot of the jet (Aurass et al., 1994; Raulin et al., 1996).

Figure 4. Coronal X-ray source above a flare of 2002 April 15, a feature interpreted by Sui & Holman (2004) as the coronal end of a reconnecting current sheet prior to its explosive eruption.

The RHESSI data, in comparison with TRACE EUV observations, confirm this association of jet, microflare, and type III burst (Figure 5). Theoretically it seems inescapable that non-thermal electrons are streaming away from the magnetic structures seen in the low corona and imaged in hard X-rays by RHESSI. Thus a sensitive analysis of the RHESSI data should show non-thermal bremsstrahlung along the jet; this would confirm at high resolution that these jet flows provide channels for electron propagation and the escape of the electromagnetic radiation they induce. The problem is not just one of sensitivity, but also dynamic range; the density in the jet itself may not exceed 10^{9} cm^{−3} (Shimojo et al., 1996) whereas the density in a thick-target footpoint

Figure 5. RHESSI hard X-ray observation (contours, 12-25 keV) with type III burst, showing the jet structure observed by TRACE.

source may reach reach 10^{14} cm^{−3}, depending upon the particle energy and transport model.

[Calculation of hard X-ray brightness in jet and estimation of requirements on sensitivity and image dynamic range for direct detection]

[Discussion of Kontar & Krucker (2005), the finding of correlations between SEP and HXR spectra]

7. Parallel motions

Lots of Sui material; Somov et al., 2005

8. Footpoint-to-looptop ratio

Melrose and Brown, 1976, MacKinnon, 1988, Fletcher and Martens, 1998

Figure 6. Place holder for a summary sketch showing when and where different kinds of coronal hard X-ray emissions occur, and what their spectral distributions are (this figure is from Wild et al., 1963. I am sure that we can do better, and at the same time avoid getting hung up in nomenclature - HH).

9. Collapsing traps

[It would be particularly interesting to hear how the collapsing-trip scenario ties in with the distributed electric fields discussed by Drake et al. (2006) and Onofri et al. (2006)

10. Conclusions

[A 3D time/space/spectrum spectrogram representation of July 23, illustrating when and how the different components appear); see the placeholder in Figure 6 for a bad beginning]

Acknowledgements

NASA NAG5-12878 (HSH)

Appendix

[If we are energetic enough, this section should contain a lot of information useful for the interpretation of hard X-ray sources in the corona. This is a list of some of the items we might think to include. In my imagination I also see us actually using some of these plots, or even the software that made them, to help analyze individual events at ISSI - HH]

Bremsstrahlung cross-section

Coulomb collisional cross-sections

Particle adiabatic motion

Magnetic models The PFSS modeling of the solar coronal magnetic field has reached a certain level of maturity and convenience (e.g., Schrijver and Derosa, 2003. For example, meter-wave images of type III radio bursts can correctly be identified with open field lines in such models (Klein, Nitta). We can probably thus begin to think of fitting specific large-scale field geometries to specific coronal hard X-ray events.

Three-dimensional projections Loughhead et al., 1983, Berton and Sakurai, 1985, Nitta, 2000

References

Acton, L. W., U. Feldman, M. E. Bruner, G. A. Doschek, T. Hirayama, H. S. Hudson, J. R. Lemen, Y. Ogawara, K. T. Strong, and S. Tsuneta: 1992, ‘The morphology of 20 X 10 exp 6 K plasma in large non-impulsive solar flares’. PASJ 44, L71–L75.

Alexander, D. and T. R. Metcalf: 2002, ‘Energy dependence of electron trapping in

a solar flare’. Solar Phys 210, 323–340. Asai, A., T. Yokoyama, M. Shimojo, and K. Shibata: 2004, ‘Downflow Motions

Associated with Impulsive Nonthermal Emissions Observed in the 2002 July 23 Solar Flare’. ApJL 605, L77–L80.

Aschwanden, M. J., L. Fletcher, T. Sakao, T. Kosugi, and H. Hudson: 1999, ‘Deconvolution of Directly Precipitating and Trap-precipitating Electrons in Solar Flare Hard X-Rays. III. Yohkoh Hard X-Ray Telescope Data Analysis’. ApJ 517, 977–989.

Aurass, H., K.-L. Klein, and P. C. H. Martens: 1994, ‘First detection of correlated electron beams and plasma jets in radio and soft x-ray data’. Solar Phys 155, 203–206.

Bai, T. and P. A. Sturrock: 1989, ‘Classification of solar flares’. ARAA 27, 421–467. Berton, R. and T. Sakurai: 1985, ‘Stereoscopic determination of the three-

dimensional geometry of coronal magnetic loops’. Solar Phys 96, 93–111. Brown, J. C.: 1971, ‘The Deduction of Energy Spectra of Non-Thermal Electrons in

Flares from the Observed Dynamic Spectra of Hard X-Ray Bursts’. Solar Phys 18, 489–+.

Cliver, E. W., B. R. Dennis, A. L. Kiplinger, S. R. Kane, D. F. Neidig, N. R. Sheeley, and M. J. Koomen: 1986, ‘Solar gradual hard X-ray bursts and associated phenomena’. ApJ 305, 920–935.

Drake, J. F., M. Swisdak, H. Che, and M. A. Shay: 2006, ‘Electron acceleration from

contracting magnetic islands during reconnection’. Nature 443, 553–556. Fárník, F., H. S. Hudson, M. Karlický, and T. Kosugi: 2003, ‘X-ray and radio

observations of the activation stages of an X-class solar flare’. A&A 399, 1159–1166.

Fletcher, L. and P. C. H. Martens: 1998, ‘A Model for Hard X-Ray Emission from

the Top of Flaring Loops’. ApJ 505, 418–431. Frost, K. J. and B. R. Dennis: 1971, ‘Evidence from Hard X-Rays for Two-Stage

Particle Acceleration in a Solar Flare’. ApJ 165, 655–+. Grigis, P. C. and A. O. Benz: 2004, ‘The spectral evolution of impulsive solar X-ray

flares’. A&A 426, 1093–1101. Hoyng, P., H. F. van Beek, and J. C. Brown: 1976, ‘High time resolution analysis

of solar hard X-ray flares observed on board the ESRO TD-1A satellite’. Solar Phys 48, 197–254.

Hudson, H. and J. Ryan: 1995, ‘High-Energy Particles In Solar Flares’. ARAA 33,

239–282. Hudson, H. S.: 1972, ‘Thick-Target Processes and White-Light Flares’. Solar Phys

24, 414–+. Hudson, H. S.: 1978, ‘A purely coronal hard X-ray event’. ApJ 224, 235–240. Hudson, H. S. and F. Fárník: 2002, ‘Spectral variations of flare hard X-rays’. In: A.

Wilson (ed.): ESA SP-506: Solar Variability: From Core to Outer Frontiers. pp. 261–264.

Hudson, H. S., T. Kosugi, N. V. Nitta, and M. Shimojo: 2001, ‘Hard X-Radiation from a Fast Coronal Ejection’. ApJL 561, L211–L214.

Hudson, H. S., R. P. Lin, and R. T. Stewart: 1982, ‘Second-stage acceleration in a

limb-occulted flare’. Solar Phys 75, 245–261. Kahler, S.: 1976, ‘Determination of the energy or pressure of a solar X-ray structure

using X-ray filtergrams from a single filter’. Solar Phys 48, 255–259. Kane, S. R., K. A. Anderson, W. D. Evans, R. W. Klebesadel, and J. Laros: 1979,

‘Observation of an impulsive solar X-ray burst from a coronal source’. ApJL 233, L151–L155.

Kane, S. R., R. W. Kreplin, M.-J. Martres, M. Pick, and I. Soru-Escaut: 1974, ‘Acceleration of electrons in absence of detectable optical flares deduced from type III radio bursts, H alpha activity and soft X-ray emission’. Solar Phys 38, 483–497.

Kiplinger, A. L.: 1995, ‘Comparative Studies of Hard X-Ray Spectral Evolution in Solar Flares with High-Energy Proton Events Observed at Earth’. ApJ 453, 973–+.

Korchak, A. A.: 1967, ‘Possible Mechanisms for Generating Hard X Rays in Solar

Flares.’. Soviet Astronomy 11, 258–+. Kosugi, T., T. Sakao, S. Masuda, K. Makishima, M. Inda, T. Murakami, Y. Ogawara,

K. Yaji, and K. Matsushita: 1992, ‘The Hard X-ray Telescope (HXT) onboard YOHKOH - Its performance and some initial results’. PASJ 44, L45–L49.

Lin, R. P.: 1970, ‘The Emission and Propagation of ˜ 40 keV Solar Flare Electrons. II: The Electron Emission Structure of Large Active Regions’. Solar Phys 15, 453–+.

Lin, R. P., S. Krucker, G. J. Hurford, D. M. Smith, H. S. Hudson, G. D. Holman, R. A. Schwartz, B. R. Dennis, G. H. Share, R. J. Murphy, A. G. Emslie, C. Johns-Krull, and N. Vilmer: 2003, ‘RHESSI Observations of Particle Acceleration and Energy Release in an Intense Solar Gamma-Ray Line Flare’. ApJL 595, L69–L76.

Lin, R. P., R. A. Schwartz, R. M. Pelling, and K. C. Hurley: 1981, ‘A new component

of hard X-rays in solar flares’. ApJL 251, L109–L114. Lin, R. P. et al..: 2002, ‘High spectral resolution measurements of a solar flare hard

X-ray burst’. Solar Phys. 210. Loughhead, R. E., J.-L. Wang, and G. Blows: 1983, ‘High-resolution photography of

the solar chromosphere. XVII Geometry of H-alpha flare loops observed on the disk’. ApJ 274, 883–899.

MacKinnon, A. L.: 1988, ‘Coulomb collisional precipitation of fast electrons in solar

flares’. A&A 194, 279–287. Masuda, S., T. Kosugi, H. Hara, S. Tsuneta, and Y. Ogawara: 1994, ‘A Loop-

Top Hard X-Ray Source in a Compact Solar Flare as Evidence for Magnetic Reconnection’. Nature 371, 495.

Melrose, D. B. and J. C. Brown: 1976, ‘Precipitation in trap models for solar hard

X-ray bursts’. MNRAS 176, 15–30. Metcalf, T. R. and D. Alexander: 1999, ‘Coronal Trapping of Energetic Flare

Particles:Yohkoh/HXT Observations’. ApJ 522, 1108–1116. Nitta, N.: 2000, ‘Flare Loop Geometry’. In: R. Ramaty and N. Mandzhavidze (eds.):

ASP Conf. Ser. 206: High Energy Solar Physics Workshop - Anticipating Hessi. pp. 258–+.

Onofri, M., H. Isliker, and L. Vlahos: 2006, ‘Stochastic Acceleration in Turbulent Electric Fields Generated by 3D Reconnection’. Physical Review Letters 96(15), 151102–+.

Parks, G. K. and J. R. Winckler: 1969, ‘Sixteen-Second Periodic Pulsations Observed in the Correlated Microwave and Energetic X-Ray Emission from a Solar Flare’. ApJ 155, L117–+.

Petrosian, V., T. Q. Donaghy, and J. M. McTiernan: 2002, ‘Loop Top Hard X-Ray

Emission in Solar Flares: Images and Statistics’. ApJ 569, 459–473. Prince, T. A., G. J. Hurford, H. S. Hudson, and C. J. Crannell: 1988, ‘Gamma-ray

and hard X-ray imaging of solar flares’. Solar Phys 118, 269–290. Qiu, J., J. Lee, and D. E. Gary: 2004, ‘Impulsive and Gradual Nonthermal Emissions

in an X-Class Flare’. ApJ 603, 335–347. Raulin, J. P., M. R. Kundu, N. Nitta, and A. Raoult: 1996, ‘Radio Continuum and

Type III Bursts Associated with Coronal X-Ray Structures’. ApJ 472, 874–+. Schrijver, C. J. and M. L. Derosa: 2003, ‘Photospheric and heliospheric magnetic

fields’. Solar Phys 212, 165–200. Somov, B. V., T. Kosugi, S. A. Bogachev, S. Masuda, and T. Sakao: 2005, ‘On

upward motions of coronal hard X-ray sources in solar flares’. Advances in Space Research 35, 1690–1699.

Warren, H. P., J. A. Bookbinder, T. G. Forbes, L. Golub, H. S. Hudson, K. Reeves, and A. Warshall: 1999, ‘TRACE and Yohkoh Observations of High-Temperature Plasma in a Two-Ribbon Limb Flare’. ApJL 527, L121–L124.

Wheatland, M. S. and D. B. Melrose: 1995, ‘Interpreting YOHKOH hard and soft

X-ray flare observations’. Solar Phys 158, 283–299. Wild, J. P., S. F. Smerd, and A. A. Weiss: 1963, ‘Solar Bursts’. ARAA 1, 291–+.

Smerd and Weiss, 1971 (March 30 event) Kontar and Krucker, in preparation

paper.tex; 9/03/2007; 14:42; p.16